OpenStudy (anonymous):

Use logarithmic differentiation to calculate the derivative. f(x) = x^(7sqrt. (x))

OpenStudy (anonymous):

start by taking the log, get \[7\sqrt{x}\ln(x)\] take the derivative using the product rule, then and the end multiply by the original function \(x^{7\sqrt{x}}\)

OpenStudy (anonymous):

ohhhhh, i forgot the product rule!! thank you :)

OpenStudy (anonymous):